# CodeScore: Evaluating Code Generation by Learning Code Execution

### v3

@@ 3.1.2 Baselines
Execution-based CEMs consist of AvgPassRatio (Hendrycks et al., 2021).
@@
    $\operatorname{AvgPassRatio}=\frac{1}{|P|}\sum_{p\in P}\frac{1}{|C_{p}|}\sum_{c\in C_{p}}\mathbb{I}\left\{\operatorname{Eval}\left(\mathbf{g}_{p},\mathcal{I}_{p,c}\right)=\mathcal{O}_{p,c}\right\},$ (10)
@@
3.1.3 Correlation Evaluation
$\tau=\frac{Concordant-Discordant}{Concordant+Discordant},$ (11)
    $r_{s}=\frac{\operatorname{cov}(\mathrm{R}(M^{1}),\mathrm{R}(M^{2}))}{\sigma_{\mathrm{R}(M^{1})}\sigma_{\mathrm{R}(M^{2})}},$ (12)
    $r_{s}=\frac{\operatorname{cov}(M^{1},M^{2})}{\sigma_{M^{1}}\sigma_{M^{2}}}.$ (13)
  • Mean Absolute Error (MAE) is a measure of errors between paired data:
    $\mathrm{MAE}=\frac{\sum_{i=1}^{N}\left|M^{1}_{i}-M^{2}_{i}\right|}{N},$ (14)
    where ||$\cdot$|| ⋅ | means the absolute-value function.
3.1.4 Implementation Details
In this paper, UniXcoder (Guo et al., 2022a) is employed as the base LLM of UniCE, which has the similar parameter size of LLMs in BERTScore (Zhang et al., 2020) and COMET (Rei et al., 2020), and larger LLMs can usually lead to better results. We train UniCE with Adam (Kingma and Ba, 2015) optimizer on a single GPU of Tesla A100-PCIe-40G. Empirically, the learning rate is set to $0.001$. The feedforward neural network of UniCE consists of 3 linear transitions with the hyperbolic tangent (Tanh) activation functions, where the corresponding output dimensions are 3,072, 1,024, and 2, respectively. The input token length is limited to 1024. To mitigate the instability of model training, we exhibit the average performance of UniCE running five times.
Table 3: Correlation comparison of functional correctness on APPS-Eval dataset.
@@ 3.1.4 Implementation Details
UniCE with $\mathcal{L}^{Uni}$ 0.1820 $\textbf{0.5275}\ \ ({\color[rgb]{1,0,0}\uparrow 40.56\%})$ $\textbf{0.7040}\ \ ({\color[rgb]{1,0,0}\uparrow 55.07\%})$ $\textbf{0.7210}\ \ ({\color[rgb]{1,0,0}\uparrow 58.87\%})$ 0.1044
@@
Table 4: Correlation comparison of functional correctness on MBPP-Eval and HE-Eval datasets.
@@
BLEU (Papineni et al., 2002) 0.1186 0.1784 1.0 $\times$ (0.87s) 0.2249 0.0678 1.0 $\times$ (0.78s)
Accuracy 0.0004 0.0299 0.1 $\times$ 0.0006 0.0367 0.1 $\times$
CodeBLEU (Feng et al., 2020) 0.1827 0.2902 5.0 $\times$ 0.3826 0.4084 6.4$\times$
CrystalBLEU (Eghbali and Pradel, 2022) 0.0295 0.1645 0.3 $\times$ 0.0158 0.2013 0.4 $\times$
@@
BERTScore (Zhang et al., 2020) 0.8842 0.1522 62.0 $\times$ 0.8862 0.0069 57.7$\times$
COMET (Rei et al., 2020) -0.5001 0.2681 69.0 $\times$ 0.0642 0.0716 58.6$\times$
CodeBERTScore (Zhou et al., 2023) 0.7863 0.2490 44.9 $\times$ 0.7917 0.2604 47.5 $\times$
@@
UniCE with $\mathcal{L}^{Ref}$ 0.2975 0.5864 17.2 $\times$ 0.3115 0.5250 30.8$\times$
UniCE with $\mathcal{L}^{Uni}$ 0.3253 0.5999 0.4055 0.6009
@@
UniCE with $\mathcal{L}^{NL}$ 0.3364 0.4492 12.6 $\times$ 0.4748 0.5217 31.0$\times$
UniCE with $\mathcal{L}^{Uni}$ 0.3327 0.5719 0.5357 0.5755
@@
UniCE with $\mathcal{L}^{Ref+NL}$ 0.2905 0.5926 20.7 $\times$ 0.3866 0.5153 33.3$\times$
UniCE with $\mathcal{L}^{Uni}$ 0.3247 $\textbf{0.6054}\ \ ({\color[rgb]{1,0,0}\uparrow 31.52\%})$ 0.4505 $\textbf{0.6048}\ \ ({\color[rgb]{1,0,0}\uparrow 19.64\%})$
@@
8 test cases per task 0.2670 0.6826 1.0k $\times$ 0.5652 0.7426 1.9k $\times$
108 test cases per task 0.2890 - 28.7k $\times$ 0.3022 - 22.1k $\times$
3.2 Experimental Results
In this section, we conduct extensive experiments to verify the effectiveness and generalization of CodeScore.
3.2.1 RQ1: Effect of CodeScore
@@
We also sought to determine the generality of CodeScore. In Table 4, we utilize CodeScore, trained on APPS-Eval, to evaluate the code in MBPP-Eval and HE-Eval with fine-tuning and zero-shot settings, respectively. It is important to note that the distributions of NL, RefCode, GenCode, and test cases across these three datasets are quite different, as evidenced by their respective statistics shown in Table 1 and Table 2. Table 4 reveals the effectiveness of CodeScore on MBPP-Eval and HE-Eval. Remarkably, CodeScore continues to achieve the best correlation compared to other match-based CEMs and LLM-based EMs in these two settings.
@@
Figure 4: The performance of Exec on MBPP-Eval and HE-Eval datasets.
3.2.2 RQ2: Effect of Exec
We also evaluate the performance of Exec on MBPP-Eval and HE-Eval datasets, as shown in Fig. 4. The experimental results indicate that Exec demonstrates remarkably high performance in terms of Precision and F1 Score. These findings suggest that Exec possesses a robust capability in identifying whether the code can be executed (when all dependencies are met), and the results also prove that using UniCE to learn code execution is effective for code evaluation.
3.2.3 RQ3: Effect of $\mathcal{L}^{Uni}$
@@
Table 5: Human evaluation for functional correctness.
@@ 3.2.4 RQ4: Human Evaluation
We randomly select 100 generated codes and evaluations scored by the five EMs mentioned above on these samples. Finally, we obtain 500 (100*5) paired generated code with reference code and NL and its EM score for human evaluation. The evaluators are computer science Ph.D. students and are not co-authors. They all have programming experience ranging from 3+ years. The 500 code snippets are divided into 10 groups, with each questionnaire containing one group. We randomly list the generated code with reference code and NL and the corresponding EM score on the questionnaire. Each group is evaluated anonymously by one evaluator, and the final score is the average of all evaluators’ scores. Evaluators are allowed to search the Internet for unfamiliar concepts.
We present the results of the human evaluation in Table 5. Remarkably, our proposed CodeScore significantly outperforms all other EMs. Relative to these, CodeScore shows an improvement of at least 54.6% in the human evaluation. All p-values are substantially less than 0.005, underscoring that these improvements are statistically significant.
@@
Figure 5: Case study. For each case, the second generated code is superior to the first one.
3.2.5 RQ5: Case Study
Fig. 5 displays a selection of generated codes and their corresponding EM scores (as per Section 3.2.4) on MBPP-Eval dataset. It becomes evident that CodeBLEU, BERTScore, and CodeBERTScore each exhibit unique issues. From these examples, we glean the following insights: 1) CodeBLEU tends to assign relatively low scores to generated code, even when the code is functionally correct. Furthermore, it appears to favor generated codes that maintain structural consistency with the reference code. For instance, even though Generated Code III.2 is functionally correct, it receives a lower CodeBLEU score than III.1, which is fundamentally incorrect. 2) Both BERTScore and CodeBERTScore have a propensity to award relatively high scores to generated code, even when the code is essentially flawed. Additionally, they often assign lower scores to better generated codes. For example, Generated Code II/III.2 has a lower BERTScore than II/III.1, and Generated Code I.2 has a lower CodeBERTScore than I.1. In contrast, CodeScore performs admirably in all of these scenarios. In summary, our proposed CodeScore aligns more closely with Ground Truth compared to other EMs. This suggests that CodeScore is more effective in estimating the functional correctness of generated code.
4 Discussion
@@
5 Related Work
In this section, we primarily discuss three types of EMs, including Match-based CEMs, Execution-based CEMs, and LLM-based EMs.
Match-based CEMs. Besides these commonly used BLEU (Papineni et al., 2002), Accuracy, and CodeBLEU (Ren et al., 2020), some niche CEMs (Popovic, 2015) are also applied to code evaluation, e.g., METEOR (Banerjee and Lavie, 2005), ROUGE (Lin, 2004), and CrystalBLEU (Eghbali and Pradel, 2022). However, these aforementioned match-based CEMs merely measure the surface-level differences in code and do not take into account the functional correctness of the generated code.
Execution-based CEMs. They attempt to handle these issues by running tests for generated code to verify its functional correctness (Kulal et al., 2019; Hendrycks et al., 2021; Hao et al., 2022). However, they come with several caveats: 1) It assumes that test cases have been given and all dependencies have been resolved. For each code generation task, supplying adequate test cases is a burden in practice, and the dependencies required vary from task to task. 2) Enormous computational overhead needs to be afforded. All generated code requires execution separately for each corresponding test case, which leads to enormous CPU and I/O overhead. 3) Execution with isolation mechanisms. The generated code could have some security risks, such as deleting files on the disk or implanting computer viruses, especially if the training data of code generation models is attacked. In a word, they are usually costly, slow, and insecure, which are often unavailable or ineffective in real-world scenarios.
LLM-based EMs. Effective evaluation of generated results is hard for both text and code generation. They likewise face the same issue of poor evaluation metrics (EMs). A recent popular trend in evaluating text generation is the design of automatic EMs based on LLMs. A part of LLM-based EMs (Rei et al., 2021; Wan et al., 2022; Rei et al., 2022) follows COMET (Rei et al., 2020) to learn high-quality human judgments of training data, which is a problem for code evaluation to obtain. Another part relies on LLM extracting token embeddings to calculate scores like BERTScore (Zhang et al., 2020), such as (Zhao et al., 2019; Sellam et al., 2020; Yuan et al., 2021; Reimers and Gurevych, 2019). A concurrent work named CodeBERTScore (Zhou et al., 2023) tries to use the same way as BERTScore with LLM pre-trained on code. However, they do not teach LLMs to learn code evaluation effectively, in other words, LLMs are still confused about how to evaluate code. Therefore, they exhibit suboptimal performance in code evaluation, as evidenced by our experimental results.
6 Conclusion and Future Work

### v4

@@ 4.1.2. Baselines
Execution-based CEM refers to AvgPassRatio (Hendrycks et al., 2021).
@@
    (10) $\operatorname{AvgPassRatio}=\frac{1}{|P|}\sum_{p\in P}\frac{1}{|C_{p}|}\sum_{c\in C_{p}}\mathbb{I}\left\{\operatorname{Eval}\left(\mathbf{g}_{p},\mathcal{I}_{p,c}\right)=\mathcal{O}_{p,c}\right\},$
@@
Table 3. Correlation comparison of functional correctness on APPS-Eval dataset.
@@
UniCE with $\mathcal{L}^{Uni}$ 0.1820 $\textbf{0.5275}\ \ ({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 40.56\%})$ $\textbf{0.7040}\ \ ({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 55.07\%})$ $\textbf{0.7210}\ \ ({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 58.87\%})$ 0.1044
@@
4.1.3. Correlation Evaluation
    (11) $\tau=\frac{Concordant-Discordant}{Concordant+Discordant},$
    (12) $r_{s}=\frac{\operatorname{cov}(\mathrm{R}(M^{1}),\mathrm{R}(M^{2}))}{\sigma_{\mathrm{R}(M^{1})}\sigma_{\mathrm{R}(M^{2})}},$
    (13) $r_{s}=\frac{\operatorname{cov}(M^{1},M^{2})}{\sigma_{M^{1}}\sigma_{M^{2}}}.$
  • Mean Absolute Error (MAE) is a measure of errors between paired data:
    (14) $\mathrm{MAE}=\frac{\sum_{i=1}^{N}\left|M^{1}_{i}-M^{2}_{i}\right|}{N},$
    where ||$\cdot$|| ⋅ | means the absolute-value function.
Table 4. Correlation comparison of functional correctness on MBPP-Eval and HE-Eval datasets.
@@ 4.1.3. Correlation Evaluation
BLEU (Papineni et al., 2002) 0.1186 0.1784 1.0 $\times$ (0.87s) 0.2436 0.0987 1.0 $\times$ (1.96s)
Accuracy 0.0004 0.0299 0.1 $\times$ 0.0011 0.0456 0.1 $\times$
CodeBLEU (Feng et al., 2020) 0.1827 0.2902 5.0 $\times$ 0.3452 0.3308 6.3 $\times$
CrystalBLEU (Eghbali and Pradel, 2022) 0.0295 0.1645 0.3 $\times$ 0.0427 0.2171 0.4 $\times$
@@
BERTScore (Zhang et al., 2020) 0.8842 0.1522 62.0 $\times$ 0.9008 0.1214 57.5$\times$
COMET (Rei et al., 2020) -0.5001 0.2681 69.0 $\times$ 0.0879 0.1437 58.2$\times$
CodeBERTScore (Zhou et al., 2023) 0.7863 0.2490 44.9 $\times$ 0.8091 0.3196 47.4 $\times$
@@
UniCE with $\mathcal{L}^{Ref}$ 0.2975 0.5864 17.2 $\times$ 0.3426 0.5671 30.2$\times$
UniCE with $\mathcal{L}^{Uni}$ 0.3253 0.5999 0.4257 0.6378
@@
UniCE with $\mathcal{L}^{NL}$ 0.3364 0.4492 12.6 $\times$ 0.4985 0.5634 30.6$\times$
UniCE with $\mathcal{L}^{Uni}$ 0.3327 0.5719 0.5624 0.6215
@@
UniCE with $\mathcal{L}^{Ref+NL}$ 0.2905 0.5926 20.7 $\times$ 0.4059 0.5965 32.9$\times$
UniCE with $\mathcal{L}^{Uni}$ 0.3247 $\textbf{0.6027}\ \ ({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 31.25\%})$ 0.4731 $\textbf{0.6597}\ \ ({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 32.89\%})$
@@
8 test cases per task 0.2670 0.6826 1.0k $\times$ 0.5994 0.6981 1.9k $\times$
108 test cases per task 0.2890 - 28.7k $\times$ 0.3695 - 21.7k $\times$
4.1.4. Implementation Details
In this paper, UniXcoder (Guo et al., 2022a) is employed as the base LLM of UniCE, which has the similar parameter size of LLMs in BERTScore (Zhang et al., 2020) and COMET (Rei et al., 2020), and larger LLMs can usually lead to better results. We format the input sequences as “[CLS] $g$ [SEP] $r$ [SEP] $n$ [SEP]”, where [CLS] and [SEP] are the special tokens in vocabulary, and we replace $g$, $r$, and $n$ with the generated code, reference code, and NL description, respectively. For the balance of three input formats during the training process, we first sample an NL along with its corresponding generated code and reference code. They are then employed to construct data in three formats: Ref-only, NL-only, and Ref&NL. Finally, these formats are combined for training UniCE. In all experiments of this paper, we train UniCE on the train set of APPS-Eval. We fine-tune UniCE on the train set of MBPP-Eval only when we specially mention it in our paper. We train UniCE with Adam (Kingma and Ba, 2015) optimizer on a single GPU of Tesla A100-PCIe-40G. Empirically, the learning rate is set to $0.001$ and the training epoch is set to 5. The feedforward neural network of UniCE consists of 3 linear transitions with the hyperbolic tangent (Tanh) activation functions, where the corresponding output dimensions are 3,072, 1,024, and 2, respectively. The input token length is limited to 1024. To mitigate the instability of model training, we exhibit the average performance of UniCE running five times.
4.2. Experimental Results
4.2.1. RQ1: Effect of CodeScore
@@
We also sought to determine the generalizability. In Table 4, we utilize CodeScore, trained on APPS-Eval, to evaluate the code in MBPP-Eval and HE-Eval with fine-tuning and zero-shot settings, respectively. It is important to note that the distributions of NL, RefCode, GenCode, and test cases across these three datasets are quite different [^9] The average length of NL, RefCode, and GenCode across these three datasets are quite different. The average length of NL in APPS-Eval is 263.8, which far exceeds MBPP-Eval (15.5) and HE-Eval (61.9). The trend of the average length of RefCode and GenCode is similar to NL. For the Average Number of test cases per task, APPS-Eval is extended from 13 to 181, while MBPP-Eval and HE-Eval are extended from 3 and 8 to 102 and 108 respectively., as evidenced by their respective statistics shown in Table 1 and Table 2. Table 4 reveals the effectiveness of CodeScore on MBPP-Eval and HE-Eval. Remarkably, CodeScore continues to achieve the best correlation compared to other match-based CEMs and LLM-based EMs in these two settings.
Table 5. Correlation comparison of functional correctness with different base models on HE-Eval dataset.
CodeScore (UniCE with $\mathcal{L}^{Uni}$)
UniXcoder 0.4731 0.4997 0.6597 0.6486 0.2179 1.00 $\times$
CodeBert 0.4809 0.4675 0.6236 0.5622 0.2344 0.98 $\times$
CodeGraphBert 0.4597 0.5073 0.6728 0.6480 0.2281 1.07 $\times$
We conduct the experiments of UniCE based on different code pre-trained models, including CodeBert, CodeGraphBert, and UniXcoder. The results of the experiments are presented in Table 5. We did not observe obvious biases when choosing different base models. One trend we observed is that the better the model’s ability to understand the code, the more accurate it is in evaluating the code.
@@
{answerbox}
Summary of RQ1:  CodeScore outperforms match-based CEMs and LLM-based EMs in terms of correlation with functional correctness, even on datasets that it was not trained on. Moreover, CodeScore operates at a speed three orders of magnitude faster than execution-based CEMs.
Figure 4. The performance of Exec on APPS-Eval, MBPP-Eval, and HE-Eval datasets.
4.2.2. RQ2: Effect of Exec
We also evaluate the performance of Exec on APPS-Eval, MBPP-Eval, and HE-Eval datasets, as shown in Fig. 4. The experimental results indicate that Exec demonstrates remarkably high performance in terms of Precision, F1 Score, and Accuracy. Through a comprehensive analysis of all datasets, we find that our approach’s performance on the APPS-Eval dataset is inferior to that on the MBPP-Eval dataset. This discrepancy is primarily due to the higher complexity and length of problems in the APPS-Eval dataset compared to those in MBPP-Eval. Furthermore, the performance on the HE-Eval dataset is the poorest, because our approach has not been trained on this dataset. Nevertheless, our approach’s performance across various metrics on the HE-Eval dataset exceeded 90% in the zero-shot setting, indicating its effective transferability to unseen datasets. These results prove that using UniCE to learn code execution is effective for code evaluation.
{answerbox}
Summary of RQ2:  The Exec component in our approach demonstrates extremely high Precision/F1 Score/Accuracy in determining whether the code can be executed when all dependencies are met.
4.2.3. RQ3: Effect of $\mathcal{L}^{Uni}$
@@
{answerbox}
Summary of RQ3:  The component $\mathcal{L}^{Uni}$ in our approach shows positive effects across different input formats.
Table 6. Human evaluation for functional correctness.
@@ 4.2.4. RQ4: Human Evaluation
The human evaluation is conducted on the Python dataset HE-Eval. We randomly select 100 samples  [^10] Considering the workload of the evaluators, we choose a moderate sample size of 100. Too many samples would exceed the evaluators’ capacity. from this dataset, each consisting of natural language descriptions, reference code, and generated code. These samples are scored using five EMs, resulting in a total of 100*5 data pairs. We invite ten computer science PhD students, each with over three years of experience in Python development, to serve as evaluators. The 500 code snippets are divided into 10 groups, with each questionnaire containing one group. We randomly list the generated code with reference code and NL and the corresponding EM score on the questionnaire. Each group is evaluated anonymously by one evaluator, and the final score is the average of all evaluators’ scores. Evaluators are allowed to search the Internet for unfamiliar concepts.
@@
Figure 5. Case study of different EMs. For each case, the second generated code is superior to the first one.
Figure 6. Case study of Exec.
We present the results of the human evaluation in Table 6. Remarkably, our proposed CodeScore significantly outperforms all other EMs. Relative to these, CodeScore shows an improvement of at least 54.6% in the human evaluation. All p-values are substantially less than 0.005  [^11] The smaller the p-value, the less likely it is that the results are due to random factors., underscoring that these improvements are statistically significant.
{answerbox}
Summary of RQ4:  Human evaluation indicates that CodeScore shows significant improvements over previous representative EMs.
4.2.5. RQ5: Case Study
Fig. 5 displays a selection of generated codes and their corresponding EM scores (as per Section 4.2.4) on MBPP-Eval dataset. It becomes evident that CodeBLEU, BERTScore, and CodeBERTScore each exhibit unique issues. From these examples, we glean the following insights: 1) CodeBLEU tends to assign relatively low scores to generated code, even when the code is functionally correct. Furthermore, it appears to favor generated codes that maintain structural consistency with the reference code. For instance, even though Generated Code III.2 is functionally correct, it receives a lower CodeBLEU score than III.1, which is fundamentally incorrect. 2) Both BERTScore and CodeBERTScore have a propensity to award relatively high scores to generated code, even when the code is essentially flawed. Additionally, they often assign lower scores to better generated codes. For example, Generated Code II/III.2 has a lower BERTScore than II/III.1, and Generated Code I.2 has a lower CodeBERTScore than I.1. In contrast, CodeScore performs admirably in all of these scenarios. Our CodeScore aligns more closely with Ground Truth compared to other EMs. Moreover, the various formats of input have little impact on CodeScore’s scorings, indicating that CodeScore can effectively make judgments based on natural language and/or reference code, adapting to different input formats.
We further examine Exec’s capabilities through a case study. We find that Exec can effectively discriminate the cases of successful and unsuccessful compilation, especially sensitive to some errors that lead to compilation failures. A representative example is shown in Figure 6, where in Generated Code 1, the code with mismatched parentheses is recognized by Exec, and in Generated Code 2, the code with multiple nested parentheses is not misidentified by Exec.
{answerbox}
Summary of RQ5:  Through case studies, we find that our approach does not have the problems faced by previous EMs and is effective in evaluating the functional correctness and compilability of generated code.
5. Threats to Validity
There are two major threats to the validity of our work. 1) Threats to external validity concern the quality of experimental datasets and the generalizability of our results. We evaluated our approach using three public code generation datasets, which are considered mainstream benchmarks in the field and have been utilized extensively in prior research (Luo et al., 2023; Li et al., 2023; Inala et al., 2022; Huang et al., 2023; Zhang et al., 2023; Wei et al., 2023). Given their widespread use, we believe that the findings derived from these datasets offer a reasonable degree of generalizability and could potentially extend to other datasets. 2) Threats to internal validity involve the impact of hyperparameters and instability characteristics of deep learning models. Deep learning models exhibit a certain sensitivity to hyperparameter settings. In our approach, we conduct a small-range grid search on hyper-parameters using a distinct validation subset. The same set of hyperparameters is consistently applied across all datasets and compared with various baselines, achieving favorable performance consistently. Even with the same hyper-parameters, deep learning models still encounter instability issues due to factors such as the random initialization of model parameters and the random shuffling of training data. Therefore, in our experiments, we run UniCE 5 times and report its average performance. For fairness, we also run other LLM-based metrics five times with their public source code and provide the average performance.
6. Discussion
@@
7. Conclusion and Future Work